OpenStudy (anonymous):

Show how you solve and find the answer to the following. If 15 m = 63, find the value of 4m.

OpenStudy (anonymous):

15 63 cross multiply and solve 4 x

OpenStudy (anonymous):

15m = 63 /:15 m = 63/15 /*4 4m = 63*4/15 4m = 252/15